In base five, the first ten counting numbers are 1, 2, 3, 4, 10, 11, 12, 13, 14, 20. What base ten number would correspond to 11112 in base 5?
a.22 b.52 c.252 d.782 e.1222

Answer by galactus(183)

To convert from some base to base 10, write the number down like so:
1
1
1
1
2
Start at the bottom and multiply by the powers of 5 starting at 5^0
1*5^4=625
1*5^3=124
1*5^2=25
1*5^1=5
2*5^0=2
Add them up and get 782